# Binding Affinity Prediction: From Conventional to Machine Learning-Based Approaches

### v1

@@ 3.1.4 Hybridization of these three methods
Efforts have been made to bring different categories of scoring functions together to improve performance, blurring the boundaries between these categories. Some physics-based methods use weight parameters derived from regressions to increase performance. Additionally, some knowledge-based potential methods add solvation and entropy terms (Liu and Wang, 2015).
@@
Multiple reviews and comparative studies have been conducted to access the convectional scoring functions. A comparative study of 16 popular conventional scoring functions by Cheng et al. (2009) indicates that no single one consistently outperforms the others in scoring, ranking, and docking. The Pearson correlation between the scores and the experimentally determined binding affinity of the most commonly used scoring functions, such as Glide (Halgren et al., 2004), AutoDock, and Dock, ranges from 0.4 to higher 0.6. This indicates a noticeable correlation and provides some justification for using docking, but it is not sufficient to be applied in virtual screening and drug discovery with high confidence. However, despite the sub-par performance in binding affinity prediction, multiple studies have found that docking software with conventional scoring functions can accurately predict the correct binding pose and conformation (60-80%) (Cheng et al., 2009; Plewczynski et al., 2011; Li et al., 2014). This accuracy is not typically achieved by descriptor-based scoring functions (Xie and Hwang, 2014).
@@
Conventional scoring functions were proposed in the early stages of research on structure-based virtual screening and are widely used in almost all commercial and academic docking software. These scoring functions are still useful and relevant to research on protein-ligand binding today. Most of them are based, to some extent, on the theory of physics and chemistry, which makes them not only reliable and explainable to some degree but also capable of improvement as the theory advances. Cavasotto and Aucar (2020) showed that incorporating a quantum mechanical semi-empirical PM7 Hamiltonian as a scoring function significantly improved the number of compounds correctly screened using molecular docking.
Another advantage of conventional scoring function is that, despite the use of statistical machine learning algorithms, their simple function form and low learning capacity allow them to perform reasonably well without training on too much data. This characteristic was extremely desirable before and during the 2000s, when there were not enough complex structures available for the training of complex learning algorithms and models.
@@
Yet despite these advantages and some recent progress, conventional scoring functions are losing momentum in both research and industry. The disadvantages of conventional scoring functions are mostly rooted in their additive functional forms (Li et al., 2015). First of all, the energy/potential terms in the additive functions are assumed to be independent with each other, which is often not the case, especially for scoring functions based on energy terms (Khamis et al., 2015). Secondly, these functions are not expressive enough for non-linearity, indicating their incapability of complex curve-fitting and conditional branching. Therefore, these scoring functions cannot take advantage of higher-quality data (Ballester et al., 2014) and the growing number of available protein-ligand complex structures. Lastly, even from a theoretical point of view, upon which the conventional methods were usually based, these scoring functions ignored many important aspects of binding, such as the implicit treatment of solvent and protein flexibility (Ballester and Mitchell, 2010a), which became sources of error that are difficult to address. All these drawbacks are related to two factors that are deeply rooted in all conventional scoring functions: (1) the underlying theory of docking is incomplete, and the scoring functions, as a way to approximate the binding affinity, are flawed in many ways (Pantsar and Poso, 2018); (2) the additive functional form neither accurately reflects the physio-chemical process of binding nor is suitable for learning from the increasing number of available protein-ligand complex structures.
3.2 Machine learning-based methods.
The high cost of conventional physics-based approaches, as implemented in tools such as Openeye, leads to growing interest in machine-learning approaches to binding affinity prediction. The scientific community has primarily pursued interaction-free and interaction-based models, as we discuss in the next two subsections. These methodologies are distinguished by whether or not they rely on physical interactions to inform predictions.
@@
Interaction-free models infer binding affinity from data without focusing on direct physical interactions. Specifically, these ML-based models consist of two separate parts, each aiming to learn representations from molecule and protein data, including SMILES, protein sequences and graphs. In this approach, the interactions between proteins and ligands are implicitly captured in the latent spaces of their embeddings, which are formed through a neural network that processes their concatenated representations.
@@
With the development of graph neural networks (GNNs), researchers began using more advanced representations. They started by representing molecules as graphs, which better captured the relationships between atoms, while still using linear sequences for proteins (Nguyen et al., 2021; Yang et al., 2022). This change allowed for more accurate modeling of molecular structures and improved the models’ predictions.
As GNNs improved, researchers began representing complex protein 3D structures as graphs too. This approach, where both proteins and molecules are represented as graphs, has greatly enhanced the models’ ability to capture the details of molecular interactions (Somnath et al., 2021; Jiang et al., 2022; Wang et al., 2023; Liu et al., 2024b). By including spatial information, these models give a more complete view of protein-ligand complexes, leading to better predictions of binding affinity.
Additionally, the growing interest in multimodal models in the ML community has led to the development of models that combine different types of data. Researchers have created multimodal models that use both 1D sequences and 3D graphs to represent proteins (Wu et al., 2023; Zhang et al., 2023; Wu et al., 2024). These models take advantage of the strengths of different data types, combining detailed sequence information from 1D sequences with spatial and relational information from 3D graphs. This combination is especially useful for capturing complex interactions that might be missed by using only one type of data.
However, even with these advancements, interaction-free models still face challenges. The inputs to these models often lack detailed interaction information, making it difficult to accurately predict protein-ligand binding affinity. Moreover, GNNs often struggle to capture essential long-range interaction information between proteins and ligands, which is crucial for predicting binding affinity accurately.
@@
(Zhang et al., 2023) Multi (Graph + Seq) Multi (Graph + Seq) Davis, KIBA
(Liu et al., 2024b) Multi (Graph + Seq) Graph PDBBind
@@
Interaction-based models make predictions based on the 3D structures of complexes and the physical interactions between proteins and ligands. These models employ only the atoms surrounding the interaction/binding pocket to build graphs for prediction. Interaction-based methods are primarily dominated by 3D voxel grids and graphs, relying on 3D CNNs and GNNs, respectively. 3D voxel grid-based methods (Jiménez et al., 2018; Li et al., 2019; Hassan-Harrirou et al., 2020; Jones et al., 2021) use a 3D voxel grid to represent the 3D structures of protein-ligand complexes as input features for CNNs. However, the 3D voxel grid representation has several limitations. First, it has high memory consumption and computational cost due to the sparsity of the voxels, as the protein structure occupies only a small part of the entire grid. Additionally, the sensitivity of 3D grids to rotation negatively impacts prediction results. In contrast, graph-based methods (Jiang et al., 2021; Li et al., 2021; Moon et al., 2022; Yang et al., 2023; Yu et al., 2023) are mostly rotation-invariant, making their graph representations more robust than grid representations. However, interaction-based models are limited to known protein-ligand complex structures, making them less useful compared to interaction-free methods when faced with unknown protein-ligand complexes.
To address these challenges, recent research has focused on developing models that can simultaneously predict high-quality protein-ligand complexes and binding affinity (Lu et al., 2022; Tan et al., 2024). These innovations hold the potential to overcome the limitations of both interaction-free and interaction-based methods by generating accurate predictions even in the absence of known complex structures.
@@
(Lu et al., 2022) Graph PDBBind
@@ 4 Evaluation
Different from most numeric regression problems, the evaluation of binding affinity prediction is much more complex than simple error assessment, which is pointed out repeatedly in multiple comparative studies and benchmarking papers (Cheng et al., 2009; Ashtawy and Mahapatra, 2012; Li et al., 2014; Khamis and Gomaa, 2015; Liu et al., 2017; Su et al., 2018). Evaluation by ranking with decoy ligands and binding poses is necessary to simulate the process of virtual screening, and hereby make sure of the model’s practical usefulness. Cheng et al. (2009) evaluated scoring functions or binding affinity predictors from three perspectives: scoring power (binding affinity prediction), docking power (binding pose prediction), and ranking power (ligands relative ranking prediction). In a more recent study by the same group, Li et al. (2014) enriched the evaluation by adding screening power (discrimination of true binders from decoys) to the set.
In the following subsections, we review the concepts, procedures, and metrics of all four aspects in the evaluation of protein-ligand binding affinity predictors.
@@ 4.1 Scoring Power
Scoring power (Su et al., 2020) refers to the ability to produce binding scores that are correlated to the experimentally measured affinities, preferably in a linear fashion. For conventional scoring functions, the scoring power could be measured with Pearson’s correlation coefficient ($R$) or with standard deviation ($SD$):
$R=\frac{\sum(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum{(x_{i}-\bar{x})}^{2}}\sqrt{\sum{(y_{i}-\bar{y})}^{2}}},~{}~{}~{}SD=\sqrt{\frac{\sum{[y_{i}-(ax_{i}+b)]}^{2}}{N-1}}$ (6)
where $x_{i}$ is the prediction for the $i$th complex in the evaluation set, $y_{i}$ is the experimental binding affinity, $a$, $b$ are linear regression factors between scores and binding affinity, and $N$ is the number of samples in the evaluation set.
However, machine learning methods can often estimate the binding affinity directly, instead of producing a score. In these cases, common regression metrics such as MAE, MSE or RMSE can also be used. Still, Pearson’s correlation coefficient ($R$) seems be the most commonly used one for its simplicity and invariance to scaling and unit. There exists other statistical indicators that reflects the linear correlation between the predictions and targets. However, these indicators often correlate with Pearson’s correlation coefficient and do not provide extra information.
Scoring power is the most essential ability of binding affinity models, which means that the evaluation of scoring power is often adequate for studies that focus solely on affinity prediction. However, in most cases, we are also curious about how the models perform in different scenarios or real-life settings, such as docking pose prediction and virtual screening. It is more than common, according to Li et al. (2014), to see a model make excellent predictions for binding affinity, but be incapable of telling the true ligand from decoys, which ultimately renders it useless in drug discovery. Hence other evaluations for docking, ranking, and screening, are required in order to assess the models in a more comprehensive way.
@@ 4.2 Docking Power
Docking power refers to the ability that a model can tell the true binding pose from the decoy poses of a given protein-ligand pair. Ideal, the complex with native binding pose should be the one with the highest score or predicted affinity, compared to all the decoys. To implement such evaluation, complexes with decoy binding poses are usually generated with molecular docking program or through molecular dynamics simulation. In the later versions of CASF (Li et al., 2014), they used multiple docking programs such as AutoDock (Huey et al., 2012), LigandFit (Venkatachalam et al., 2003), GOLD (Verdonk et al., 2003), Surflex (Jain, 2003), and FlexX (Schellhammer and Rarey, 2004) for decoy pose generation, to minimize the bias in conformation sampling. The native/true pose is explicitly added into the decoy set so that there is at least one correct binding pose in CASF, according to Cheng et al. (2009). After the scores/affinities for all decoy poses are generated, we compare the pose with the highest score/affinity against the true pose using RSMD, which is calculated with:
@@
where $(x_{i},y_{i},z_{i})$ and $(x^{\prime}_{i},y^{\prime}_{i},z^{\prime}_{i})$ are the coordinates of the $i$th atom in the true and predicted decoy poses separately.
Lower RMSD is more desirable because it means that the predicted pose is closer to the native one. Therefore, we consider a model successfully predicted the docking pose for a protein-ligand pair if the RMSD between the true and predicted poses is below a given cutoff. In CASF, the docking power of a model is evaluated by the success rate under different cutoffs ($1.0$, $2.0$, and $3.0$) for RMSD.
One can imagine there are multiple ways to evaluate the docking power. For example, the RMSD between the native and predicted poses can be directly used for assessment. Alternatively, we can simply use the accuracy for predicting the complex in native binding pose with the highest score/affinity. However, due to the resolution and different structure measurement techniques, the native poses obtained experimentally might not be the truly optimized ones. Also, we have to take the flexibility of proteins into account. Therefore, using a static pose directly without a cutoff might cause misleading evaluation results.
Aside from the aforementioned evaluation issue, RMSD, as a metric for distance deviation, has been questioned in multiple independent studies for its drawbacks, such as sensitiveness to the atom ordering, and unawareness of the flexibility of protein structures. Li et al. (2014) used RMSDPM which matches atom pairs between two poses using atom types instead of the atom IDs. Damm and Carlson (2006) proposed wRMSD, a weighted alternative to RMSD that takes into account the flexibility of proteins, which was then recommended by Carlson (2016) as a better alternative to naive RMSD.

### v2

@@ 3.1.4 Hybrid methods
Efforts have been made to bring different categories of scoring functions together to improve performance, blurring the boundaries between these categories. Some physics-based methods use weight parameters derived from regressions to increase performance. Furthermore, some knowledge-based potential methods add solvation and entropy terms (Liu and Wang, 2015).
@@
Multiple reviews and comparative studies have been conducted to assess the conventional scoring functions. A comparative study of 16 popular conventional scoring functions Cheng et al. (2009) indicates that no single one consistently outperforms the others in scoring, ranking, and docking. The Pearson correlation between predicted scores from the most commonly used scoring functions (e.g., Glide (Halgren et al., 2004), AutoDock, Dock) and the experimentally-determined binding affinity ranges from 0.4 to 0.6. This evidence provides some justification for the use of docking, but it is not sufficient to be applied in virtual screening and drug discovery with high confidence. Despite the subpar performance in the prediction of binding affinity, studies have found that binding software with conventional scoring functions can accurately predict the correct binding pose and conformation (60-80%) (Cheng et al., 2009; Plewczynski et al., 2011; Li et al., 2014a). This precision is typically not achieved by descriptor-based scoring functions (Xie and Hwang, 2014).
@@
Conventional scoring functions were proposed in the early stages of research on structure-based virtual screening and are widely used in almost all commercial and academic docking software. These scoring functions are still useful and relevant to research on protein-ligand binding today. Most of them are based, to some extent, on the theory of physics and chemistry, which makes them not only reliable and explainable to some degree but also capable of improvement as the theory advances. Cavasotto and Aucar (2020) shows that incorporating a PM7 semiempirical quantum mechanical method as a scoring function significantly improves the number of compounds correctly screened using molecular docking. Another advantage of conventional scoring function is that—despite the use of statistical machine learning algorithms—their simple function form and low learning capacity allow them to perform reasonably well without training on too much data. This characteristic was extremely desirable before and during the 2000s, when there were not enough complex structures available for the training of complex learning algorithms and models.
@@
Yet despite these advantages and some recent progress, conventional scoring functions are losing momentum in both research and industry. The disadvantages of conventional scoring functions are mostly rooted in their additive functional forms (Li et al., 2015). First, the energy/potential terms in the additive functions are assumed to be independent with each other, which is often not the case, especially for scoring functions based on energy terms (Khamis et al., 2015). Secondly, these functions are not expressive enough for non-linearity, indicating their incapability of complex curve-fitting and conditional branching. As such, these scoring functions cannot take advantage of the growing number of available higher-quality protein-ligand complex structures data (Ballester et al., 2014). Lastly, even from a theoretical point of view, upon which the conventional methods were usually based, these scoring functions ignored many important aspects of binding, such as the implicit treatment of solvent and protein flexibility (Ballester and Mitchell, 2010a), which became sources of error that are difficult to address. All these drawbacks are related to two factors that are deeply rooted in all conventional scoring functions: (1) the underlying theory of docking is incomplete, and the scoring functions, as a way to approximate the binding affinity, are flawed (Pantsar and Poso, 2018); (2) the additive functional form neither accurately reflects the physio-chemical process of binding nor is suitable for learning from the increasing number of available protein-ligand complex structures.
Conventional scoring functions provided the first generation of computational predictors, balancing efficiency and interpretability. Yet, as structural databases expanded and computational power increased, their rigid formulations proved inadequate for capturing the complexity of protein–ligand interactions, thereby catalyzing the rise of machine learning–based methodologies.
3.2 Machine Learning-based Methods
The prohibitive cost and scalability limits of conventional physics-based approaches (e.g., OpenEye) have motivated increasing reliance on machine learning for binding affinity prediction. Within this paradigm, two complementary directions have emerged: interaction-free models, which are primarily representation-driven and infer affinities from learned embeddings of proteins and ligands without explicit structural interactions, and interaction-based models, which are more physics-informed and ground their predictions in the spatial and chemical features of binding pockets. In addition, during prediction, a variety of representations are employed to capture the structural and chemical characteristics of proteins and molecules. Sequence-based approaches model proteins as amino acid sequences and molecules as SMILES strings, emphasizing their primary structural information. Graph representations encode proteins by treating atoms as nodes with edges formed via k-nearest neighbors from 3D coordinates, while molecules are represented with atoms as nodes and chemical bonds as edges, capturing topological relationships. Voxel-based methods discretize protein structures into 3D grids that encode the spatial arrangement of binding sites for convolutional modeling. Finally, point cloud representations describe proteins as sets of atomic coordinates in 3D space, facilitating direct geometric reasoning over spatial distributions. This dichotomy reflects a broader philosophical divide between abstract representation learning and explicit interaction modeling, framing much of the recent methodological innovation in the field.
@@ 3.2.1 Interaction-free approach
Interaction-free models infer binding affinity from data without focusing on direct physical interactions. Specifically, these ML-based models consist of two separate parts, each aiming to learn representations from molecule and protein data, including SMILES, protein sequences and graphs. The interactions between proteins and ligands are implicitly captured in the latent spaces of their embeddings, which are formed through a neural network that processes their concatenated representations.
@@
With the development of graph neural networks (GNNs), researchers began using more advanced representations. They started by representing molecules as 2D molecular graphs, where atoms are nodes and bonds are edges, thereby capturing the topological relationships between atoms. Proteins, however, are still typically represented as linear amino acid sequences (Nguyen et al., 2021; Yang et al., 2022). This change allowed for more accurate modeling of molecular structures and improved the models’ predictions.
As GNNs improved, researchers began representing complex protein 3D structures as graphs too. This approach, where both proteins and molecules are represented as graphs, has greatly enhanced the models’ ability to capture the details of molecular interactions (Somnath et al., 2021; Jiang et al., 2022; Wang et al., 2023). By including spatial information, these models enhances representation fidelity by explicitly modeling 3D topological and spatial constraints of protein-ligand complexes, leading to better predictions of binding affinity.
Additionally, the growing interest in multimodal models has led to the development of multimodal models that combine both 1D sequences and 3D graphs to represent proteins (Wu et al., 2023; Zhang et al., 2023; Wu et al., 2024; Liu et al., 2025). These models take advantage of the strengths of different data types, combining detailed sequence information from 1D sequences with spatial and relational information from 3D graphs. This combination is especially useful for capturing complex interactions that might be missed by using only one type of data. Notably, large-scale pretrained protein language models have been increasingly adopted (Wu et al., 2023, 2024; Liu et al., 2025). These methods rely on pretrained protein language models such as ESM (Lin et al., 2023) to derive embeddings from raw sequences. By leveraging large-scale pretraining on millions of protein sequences, these embeddings capture generalizable biochemical and evolutionary features, which can then be integrated with structural representations. Such pretraining strategies have been shown to improve generalization to unseen proteins, underscoring the scalability and adaptability of interaction-free approaches within multimodal settings.
@@
(Liu et al., 2025) Multi (Graph + Seq) Graph PDBBind
@@
However, even with these advancements, interaction-free models face challenges. The inputs to these models often lack detailed interaction information, making it difficult to accurately predict protein-ligand binding affinity. Moreover, GNNs often struggle to capture essential long-range interaction information between proteins and ligands, which is crucial for predicting binding affinity accurately.
@@ 3.2.2 Interaction-based approach
Interaction-based models make predictions based on the 3D structures of complexes and the physical interactions between proteins and ligands. These models employ only the atoms surrounding the interaction/binding pocket to build graphs for prediction.
Interaction-based methods are primarily dominated by 3D voxel grids and graphs, relying on 3D CNNs and GNNs, respectively. 3D voxel grid-based methods (Jiménez et al., 2018; Li et al., 2019; Hassan-Harrirou et al., 2020; Jones et al., 2021) use a 3D voxel grid to represent the 3D structures of protein-ligand complexes as input features for CNNs. However, the 3D voxel grid representation has several limitations. First, it has high memory consumption and computational cost due to the sparsity of the voxels, as the protein structure occupies only a small part of the entire grid. Additionally, the sensitivity of 3D grids to rotation negatively impacts prediction results. In contrast, graph-based methods (Jiang et al., 2021; Li et al., 2021; Moon et al., 2022; Yang et al., 2023; Yu et al., 2023) are mostly rotation-invariant, making their graph representations more robust than grid representations. Still, interaction-based models are limited to known protein-ligand complex structures, making them less useful compared to interaction-free methods when faced with unknown protein-ligand complexes.
To address these challenges, recent research has begun to develop models capable of jointly predicting high-quality protein–ligand complex structures and their binding affinities (Lu et al., 2022; Tan et al., 2024). Such approaches aim to bridge the gap between interaction-free and interaction-based paradigms, offering the potential to generate reliable predictions even in the absence of experimentally resolved complexes. Although there are works combining complex prediction and binding affinity prediction, most recent structure prediction models (Jumper et al., 2021; Krishna et al., 2024; Abramson et al., 2024; Liu et al., 2024a) while able to predict complex structures, are not designed to predict binding affinities. In contrast, the most recent Boltz2 (Passaro et al., 2025) framework supports both binding affinity prediction and complex structure prediction, highlighting a step forward in unifying these tasks.
Beyond the methodological advances, another emerging direction is the integration of dynamics and flexibility into complex representations. Most existing approaches treat protein–ligand interactions as static snapshots, neglecting the conformational variability that can critically influence binding affinity. Recent works have begun to explore incorporating molecular dynamics simulations (Min et al., 2024; Passaro et al., 2025) or diffusion-based generative frameworks (Guan et al., 2023; team et al., 2024; Wohlwend et al., 2025; Team et al., 2025; Lin et al., 2025; Passaro et al., 2025) to model the ensemble nature of protein–ligand complexes. By explicitly accounting for conformational changes and induced fit effects, these approaches aim to better approximate real binding processes and improve generalization to diverse biological systems. Such dynamic-aware methods represent a promising avenue for bridging the gap between computational predictions and experimentally observed binding behaviors.
@@
(Guan et al., 2023) Graph PDBBind, CrossDocked2020
@@
(Passaro et al., 2025) Point cloud PubChem, ChEMBL, BindingDB
@@
Advances in both interaction-free and interaction-based machine learning approaches signal a decisive shift beyond the limitations of conventional scoring functions, with predictive improvements emerging alongside broader transformations in drug discovery, such as the FDA’s recent phase-out of mandatory animal testing. Within this evolving landscape, AI-driven in silico frameworks—most notably AI Virtual Cells (AIVCs)—are poised to redefine binding affinity prediction by situating molecular interactions within dynamic, multi-omic, and cell-type–specific contexts. Progress in affinity prediction will directly enhance the fidelity of AIVCs, while advances in AIVCs will reciprocally provide richer system-level environments to refine and validate predictive models, creating a mutually reinforcing cycle that could ultimately enable simulations of temporal dynamics, system-level specificity, and more personalized therapeutic outcomes. Yet, critical gaps remain, particularly in addressing conformational flexibility, dataset bias, and the integration of multi-omics information—challenges we explore further in Section 5.
@@ 4 Evaluation
Different from most numeric regression problems, the evaluation of binding affinity prediction is much more complex than simple error assessment, which has been pointed out repeatedly in multiple comparative studies and benchmarking papers (Cheng et al., 2009; Ashtawy and Mahapatra, 2012; Li et al., 2014a; Khamis and Gomaa, 2015; Liu et al., 2017; Su et al., 2018). Evaluation by ranking with decoy ligands and binding poses is necessary to simulate the process of virtual screening, ensuring the model’s practical usefulness. Cheng et al. (2009) evaluated scoring functions or binding affinity predictors from three perspectives: scoring power (binding affinity prediction), docking power (binding pose prediction), and ranking power (ligands relative ranking prediction). In a more recent study by the same group, Li et al. (2014a) enriched the evaluation by adding screening power (discrimination of true binders from decoys) to the set.
@@ 4.1 Scoring Power
Scoring power (Su et al., 2020) refers to the ability to produce binding scores that are correlated with the experimentally-measured affinities, preferably in a linear fashion. For conventional scoring functions, the scoring power could be measured with Pearson’s correlation coefficient ($R$) or standard deviation ($\mathrm{SD}$):
$R=\frac{\sum(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum{(x_{i}-\bar{x})}^{2}}\sqrt{\sum{(y_{i}-\bar{y})}^{2}}},\qquad\mathrm{SD}=\sqrt{\frac{\sum{[y_{i}-(ax_{i}+b)]}^{2}}{N-1}}$ (21)
where $x_{i}$ is the prediction for the $i$th complex in the evaluation set, $y_{i}$ is the experimental binding affinity, $a$ and $b$ are linear regression factors between the predicted scores and binding affinity, and $N$ is the number of samples in the evaluation set.
However, machine learning methods can often estimate the binding affinity directly instead of producing a score. In these cases, common regression metrics such as Mean Absolute Error (MAE), Mean Squared Error (MSE) or Root Mean Squared Error (RMSE) can also be used. Still, Pearson’s correlation coefficient ($R$) seems be the most commonly used one for its simplicity and invariance to scaling and unit. There exists other statistical indicators that reflect the linear correlation between the predictions and targets. However, these indicators often correlate with Pearson’s correlation coefficient and do not provide extra information.
Scoring power is the most essential ability of binding affinity models, which means that the evaluation of scoring power is often adequate for studies that focus solely on affinity prediction. However, in most cases, we are also curious about the models’ performance in different settings, such as docking pose prediction and virtual screening. It is more than common for a model to make excellent predictions for binding affinity, but be incapable of differentiating between the true ligand from decoys, ultimately rendering the model useless in drug discovery (Li et al., 2014a). As such, other evaluations for docking, ranking, and screening are required to assess the models in a more comprehensive way.
@@ 4.2 Docking Power
Docking power refers to the ability of a model to differentiate between the true binding pose from the decoy poses of a given protein-ligand pair. Ideally, the complex with the native binding pose should have the highest score or predicted affinity compared to the decoys. To implement such evaluation, complexes with decoy binding poses are usually generated with a molecular docking program or through molecular dynamics simulation. The later versions of CASF (Li et al., 2014a) used multiple docking programs (e.g., AutoDock (Huey et al., 2012), LigandFit (Venkatachalam et al., 2003), GOLD (Verdonk et al., 2003), Surflex (Jain, 2003), FlexX (Schellhammer and Rarey, 2004)) for decoy pose generation to minimize the bias in conformation sampling. The native/true pose is explicitly added into the decoy set so that there is at least one correct binding pose in CASF (Cheng et al., 2009). After the scores/affinities for the decoy poses are generated, poses with the highest score/affinity are compared against the true pose using RSMD, which is calculated with:
@@
where $(x_{i},y_{i},z_{i})$ and $(x^{\prime}_{i},y^{\prime}_{i},z^{\prime}_{i})$ are the coordinates of the $i$th atom in the true and predicted decoy poses. Lower RMSD indicates that the predicted pose is closer to the native one.
Using a static pose directly without a cutoff might cause misleading evaluation results: (1) due to the resolution and different structure measurement techniques, the native poses obtained experimentally might not be truly optimized; (2) we should consider the flexibility of proteins. So, a model is considered to have successfully predicted the docking pose for a protein-ligand pair if the RMSD between the true and predicted poses is below a given cutoff. In CASF, the docking power of a model is evaluated by the success rate under different cutoffs ($1.0$, $2.0$, and $3.0$ Å) for RMSD.
Still, using RMSD as a metric for distance deviation has drawbacks, such as sensitiveness to the atom ordering and unawareness of the flexibility of protein structures. Li et al. (2014a) used RMSDPM, which matches atom pairs between two poses using atom types instead of the atom IDs. Damm and Carlson (2006) proposed wRMSD, a weighted alternative to RMSD that takes into account the flexibility of proteins, which was then recommended by Carlson (2016) as a better alternative to naive RMSD.